OpenStudy (rissa):

Find the constant of variation k for the direct variation. 4x = -6y (A.) k = -2/3 (B.) k = 2/3 (C.) k = -3/2 (D.) k = -6

OpenStudy (dumbcow):

y = kx -> y = -4/6x = -2/3x therefore, k = -2/3